OpenStudy (anonymous):

A wheel is shaped like a circle with a diameter of 35 cm. To the nearest centimeter, what is the circumference of the wheel? A. 220 cm B. 110 cm C. 55 cm D. 70 cm please give me the best answer.

OpenStudy (anonymous):

circumference=2*3.14*radius. radius=35/2=17.5

OpenStudy (anonymous):

circumference=2*3.14*17.5=110